# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

### v2

@@ 2.3. Architecture Engineering
Different from above methods of generating modality or representations of modalities, some researchers adjust the model architecture to adapt to the missing-modality cases. We divide them into four categories according to their core contributions in dealing with missing modalities: attention-based methods, distillation-based methods, graph learning-based methods, and multimodal large language models (MLLMs).
(a) Intra-Modality Attention Method
(b) Inter-Modality Attention Method
@@
Figure 9. General idea of two typical attention fusion methods. We set modality 2 as the missing modality. In (a), the missing modality 2 in intra-modality attention (“Intra-Atten” in figure) can be skipped. Since each Intra-Atten of different modalities shares the fusion representation (like learnable query in (Ge et al., 2023) or bottleneck tokens in (Nagrani et al., 2021)), after calculating them one by one, the shared learnable representation can be regarded as the fusion of representations of available modalities. In (b), “Missing Modality Mask” is the custom attention mask that is generated according to the missing modality 2 and can help inter-modality attention (“Inter-Atten” in figure) ignore the missing modality 2. By setting the tokens of modality 2 to zero or negative infinity through masking, we can force the attention mechanism to ignore the missing modality 2.
2.3.1. Attention Based Method
@@
In the self-attention mechanism (Vaswani et al., 2017), each input is linearly transformed to generate Query, Key, and Value vectors. Attention weights are computed by multiplying the query of each element with the keys of others, followed by scaling and softmax to ensure the weights sum to 1. Finally, a weighted sum of the values generates the output. We classify attention-based MLMM methods into two categories: attention fusion methods, which integrate multimodal information (e.g., text, vision, audio) and are suitable for various networks; and transformer-based methods, which stack attention layers to handle large-scale data with global information capture and parallelization.
@@
Attention Fusion Based Methods: Due to the powerful ability of attention to capture key features and its plug-and-play benefits, many methods have adopted attention mechanisms. We categorize them into two types: intra- and inter-modality attention methods. Intra-modality attention methods compute attention for each modality independently before fusing them, as shown in Figure 9(a). This approach focuses on relationships within a single modality (e.g., text, image, or audio), and the fusion between modalities is achieved by learnable fusion representations. For instance, in 3D detection for autonomous vehicles, the BEV-Evolving Decoder (Ge et al., 2023) handles sensor failures by sharing same BEV-query with each modality-specific attention modules, allowing fusion of any number of modalities. Similarly, in clinical diagnosis, Lee et al. (Lee et al., 2023a) proposed modality-aware attention to perform intra-modality attention and predict decisions by fusing bottleneck fusion tokens (Nagrani et al., 2021), addressing learning with missing modalities. Inter-modality attention methods, often based on masked attention, treat missing modality features as masked vectors to better capture dependencies across modalities, as illustrated in Figure 9(b). Unlike conventional cross-modal attention, masked attention models share the same parameters across all embeddings, allowing flexible handling of missing modalities. For example, Qian et al.(Qian and Wang, 2023) designed an attention mask matrix to ignore missing modalities, improving model robustness. Similarly, DrFuse(Yao et al., 2024) decouples modalities into specific and shared representations, using the shared ones to replace missing modalities, with a custom mask matrix to help model ignore the specific representations of the missing modality.
@@
Figure 10. Description of general idea of joint representation learning methods. Modality 2 is set as the missing modality. The missing modality 2 tokens in can be replaced by learnable tokens or specific tokens. The modality encoders (red dotted rectangle) can be replaced by linear projection layers in some methods.
@@
Transformer Based Methods: We divide transformer-based methods into two methods: joint representation learning (JRL) and parameter efficient learning (PEL) according to full parameter training and a small amount of parameter fine-tuning. By utilizing stacked multimodal transformer layers with attention, the multimodal transformer can learn joint representations from any number of modality tokens (Figure 10). Gong et al. (Gong et al., 2023) introduced an egocentric multimodal task, proposing a transformer-based fusion module with a flexible number of modality tokens and a cross-modal contrast alignment loss to map features into a common space. Similarly, Mordacq et al. (Mordacq et al., 2024) leveraged a Masked Multimodal Transformer, treating missing modalities as masked tokens for robust joint representation learning. Ma et al. (Ma et al., 2022) proposed an optimal fusion strategy search strategy to enhance the performance of multimodal transformers in handling missing modalities. Radosavovic et al.(Radosavovic et al., 2024) introduced mask tokens [M] for an autoregressive transformer to manage missing modalities, successfully deploying the model in real-world scenarios. With the rise of pre-trained transformer models, PEL methods have been developed to fine-tune these models by training few parameters, leveraging their generalization abilities for various tasks. Given their focus on pre-trained models and limited parameter updates, we categorize PEL methods separately, despite some overlap with above JRL methods. Two common fine-tuning methods for pre-trained models are prompt and adapter tuning. Initially used in natural language processing, prompt tuning optimizes input prompts while keeping model parameters fixed. It has since been extended to multimodal models, optimizing multimodal prompt embedding to handle diverse inputs like text and images. Jang et al.(Jang et al., 2024) introduced modality-specific prompts to address limitations in earlier methods, which merge when all modalities are present and allow the model to update all learnable prompts during training. Liu et al.(Liu et al., 2024b) further improved this by proposing Fourier Prompt, which uses fast Fourier transform to encode global spectral information into learnable prompt tokens which can be used to supplement missing-modality features, enabling cross-attention with feature tokens to address missing modalities. Adapter tuning, on the other hand, involves inserting lightweight adapter layers into pre-trained models to adapt to new tasks without modifying the original parameters. Qiu et al. (Qiu et al., 2023) proposed a method that uses a classifier to identify missing-modality cases and the intermediate features as the missing-modality prompt to cooperate with a lightweight Adapter to address missing modality problems for brain tumor segmentation.
@@
Although attention-based fusion mechanisms in above methods can effectively help deal with the missing modality problem in any framework, none of them cares about the missing modality that may contain important information required for prediction. In the transformer-based methods, although the PEL method has a very fast fine-tuning speed and requires very few computing resources, its performance is still not comparable to that of JRL. But the JRL method is limited by a large amount of computing resources and requires relatively large datasets to achieve good performance.
@@
(a) Mean Teacher Distillation Method
(b) Self Distillation Method
@@
Figure 12. Description of two process-based distillation methods. (a) “EMA” means exponential moving average. To demonstrate the general architecture of Mean Teacher Distillation, we set the loss on the logits representation. (b) Self-distillation methods involve models acting as both the teachers and students, where most of them uses their own predictions from each branch to refine their representations during training.
2.3.2. Distillation Based Method
Knowledge distillation transfers knowledge from a teacher model to a student model (Hinton et al., 2015). The teacher model, with access to more information, helps the student reconstruct missing modalities. Below, we categorize two types of distillation methods for addressing this problem. First is the representation-based distillation, this method transfers rich representations from the teacher model to help the student capture and reconstruct missing modality features. We classify approaches based on whether they use decision-level or intermediate features. Figure 11 illustrates the procedures. Response distillation methods focus on transferring the teacher model’s logits to the student, helping it mimic probability distributions. Wang et al. (Wang et al., 2020b) trained modality-specific teachers for missing modalities, then used their soft labels to guide a multimodal student. Hafner and Ban (Hafner and Ban, 2023) employed logits from a teacher model trained with optical data to supervise a reconstruction network for approximating missing optical features from radar data. Pramit et al. (Saha et al., 2024) proposed a Modality-Aware Distillation, leveraging both global and local teacher knowledge in a federated learning setting to handle missing modalities. Orzikulova et al. (Orzikulova et al., 2024) used logits for privacy-preserving knowledge transfer in missing modality scenarios. Intermediate distillation methods align intermediate features between teacher and student models. Shen et al. (Shen and Gao, 2019) used Domain Adversarial Similarity Loss to align intermediate layers of teacher and student, improving segmentation in missing modality settings. Zhang et al.(Zhang et al., 2023c) applied intermediate distillation in endometriosis diagnosis by distilling features from a TVUS-trained teacher to a student using MRI data. Then is the process-based method, which focuses on overall distillation strategies, like Mean Teacher Distillation (MTD) (Tarvainen and Valpola, 2017) and self-distillation(Zhang et al., 2019). These methods emphasize procedural learning over direct representation transfer. MTD enhances stability by using the exponential moving average of the student model’s parameters as a teacher (Figure 12(a)). Chen et al. (Chen et al., 2024a) applied this to missing modality sentiment analysis, treating missing samples as augmented data. Li et al.(Li et al., 2022a) used MTD for Lidar-Radar segmentation, improving robustness against missing modalities. Self-distillation helps a model improve by learning from its own soft representations (Figure 12(b)). Hu et al. proposed the ShaSpec (Wang et al., 2023a), which utilizes distillation between modality-shared branches and modality-specific branches of available modalities to help improve model performance. Based on ShaSpec, researchers proposed Meta-learned Cross-modal Knowledge Distillation (Wang et al., 2024b) to further weigh the importance of different modalities to improve performance. Ramazanova et al. (Ramazanova et al., 2024) used mutual information and self-distillation for egocentric video tasks, making predictions invariant to missing modalities. Shi et al. proposed PASSION (Shi et al., 2024), a self-distillation method designed to use the multi-modal branch to help other uni-modal branches of avaiable modalities to improve multimodal medical segmentation performance with missing modality. Hybrid distillation methods combine various distillation approaches to improve student performance. For medical segmentation, Yang et al. (Yang et al., 2022) distilled teacher model knowledge at every decoder layer, outperforming ACNet(Wang et al., 2021). Wang et al. (Wang et al., 2023d) introduced ProtoKD, which captures inter-class feature relationships for improved segmentation under missing modality conditions. For sentiment analysis, CorrKD (Li et al., 2024a) leverages contrastive distillation and prototype learning to enhance performance in uncertain modality scenarios. Aforementioned methods effectively address the missing modality problem and achieve good generalization during testing. However, except for some intermediate and self-distillation methods, most assume the complete dataset is available for training (means teacher can access full-modality information during training), with missing modalities encountered only during testing. As a result, most are unsuitable for handling incomplete training datasets.
@@ 2.3.2. Distillation Based Method
Figure 13. Illustration of graph fusion methods. We set modality 2 (white circle) as the missing modalities. Features from modality 1 and 3 can be aggregated via this graph fusion method and fused features maintains the same dimension.
2.3.3. Graph Learning Based Method
@@
Graph-learning based methods leverage relationships between nodes and edges in graph-structured data for representation learning and prediction. We categorize approaches for addressing the missing modality problem into two main types: graph fusion and graph neural network (GNN) methods. Graph fusion methods integrate multimodal data using a graph structure (Figure 13), making them adaptable to various networks. For example, Angelou et al.(Angelou et al., 2019) proposed a method mapping each modality to a common space using graph techniques to preserve distances and internal structures. Chen et al.(Chen and Zhang, 2020) introduced HGMF, which builds complex relational networks using hyper edges that dynamically connect available modalities. Zhao et al.(Zhao et al., 2022) developed Modality-Adaptive Feature Interaction for Brain Tumor Segmentation, adjusting feature interactions across modalities based on binary existence codes. More recently, Yang et al.(Yang et al., 2023b) proposed a Graph Attention Based Fusion Block to adaptively fuse multimodal features, using attention-based message passing to share information between modalities. These fusion methods can be flexibly inserted into any network for integrating multiple modalities.
(a) Individual GNN Method
(b) Unified GNN Method
@@
Figure 14. Illustration of graph learning based methods. We set modality 2 as the missing modalities. (a) Each modality first passes through DNN or builds its graph, and then they are uniformly processed using GNNs. (b) This method first complete the graph from missing-modality samples via approaches like FeatProp (Malitesta et al., 2024), then process it by Graph Neural Networks (GNN).
@@
GNN methods directly encode multimodal information into a graph structure, with GNNs used to learn and fuse this information. Early approaches (Zhao et al., 2016; Wang et al., 2015) employed Laplacian graphs to connect complete and incomplete multimodal samples. Individual GNN methods (Figure 14(a)), such as DESAlign (Wang et al., 2024c), extract features using neural networks or GNNs and fuse them for prediction. Unified GNN methods ( Figure 14(b)) first complete the graph and then use GNNs for prediction, such as in Zhang et al.’s M3Care (Zhang et al., 2022), which uses adaptive weights to integrate information from similar patients. Lian et al.(Lian et al., 2023) proposed the Graph Completion Network, which reconstructs missing modalities by mapping features back to the input space. In recommendation systems, FeatProp(Malitesta et al., 2024) propagates known multimodal features to infer missing ones, while MUSE (Wu et al., 2024) represents patient-modality relationships as a bipartite graph and learns unified patient representations. In knowledge graphs, Chen et al. (Chen et al., 2023a) introduced Entity-Level Modality Alignment to dynamically assign lower weights to missing or uncertain modal information, reducing the risk of misguidance in learning. Above methods can leverage the graph structure to better capture relationships both within/between modalities and samples. However, these methods tend to have lower efficiency, scalability, and higher development complexity compared to other kind of approaches.
2.3.4. Multimodal Large Language Model

### v3

@@ 4.1. Architecture-Focused Models
Different from above methods of handling missing modality problems at the modality or representation level, many researchers adjust the model training or testing architecture to adapt to the missing-modality cases. We divide them into four categories according to their core contributions in dealing with missing modalities: attention-based methods, distillation-based methods, graph learning-based methods, and multimodal large language models.
(a) Intra-Modality Attention Methods
(b) Inter-Modality Attention Methods
@@
Figure 9. Two typical attention fusion methods, assuming modality 2 as the missing modality. (a) The intra-modality attention (“Intra-Atten” in figure) of missing modality 2 can be skipped. Since each Intra-Atten of different modalities shares the fusion representation (like learnable query in (Ge et al., 2023) or bottleneck tokens in (Nagrani et al., 2021)), after calculating them one by one, the output representation can be regarded as the fusion of available modalities. (b) “Missing Modality Mask” is the custom mask that is generated according to the missing modality 2 and can help inter-modality attention (“Inter-Atten” in figure) ignore the missing modality 2. By setting the tokens of modality 2 to zero or negative infinity through masking, we can force the attention mechanism to ignore the missing modality 2.
4.1.1. Attention Based Methods
@@
In the self-attention mechanism (Vaswani et al., 2017), each input is linearly transformed to generate Query, Key, and Value vectors. Attention weights are computed by multiplying the query of each element with the keys of others, followed by scaling and softmax to ensure the weights sum to 1. Finally, a weighted sum of the values generates the output. We classify attention-based MLMM methods into two categories. (1) Attention fusion methods put attention on modality fusion, integrating multimodal information, which do not rely on any specific model type and can be suitable for various model types as its input and output dimensions are the same. (2) Others are transformer-based methods, which stack attention layers to handle large-scale data with global information capture and parallelization. We provide more details of these two methods below.
@@
Attention Fusion Methods have the powerful ability to capture key features and can be seen as plug-and-play modules. We categorize them into two types: intra- and inter-modality attention methods. Intra-Modality Attention Methods compute attention for each modality independently before fusing them, as shown in Figure 9(a). This approach focuses on relationships within a single modality, and the fusion between modalities is achieved by sharing partial information. For instance, in 3D detection, the BEV-Evolving Decoder (Ge et al., 2023) handles sensor failures by sharing same BEV-query with each modality-specific attention modules, allowing fusion of any number of modalities. Similarly, in clinical diagnosis, Lee et al. (Lee et al., 2023a) proposed modality-aware attention to perform intra-modality attention and predict decisions by using shared bottleneck fusion tokens (Nagrani et al., 2021). Inter-Modality Attention Methods, often based on masked attention, treat missing modality features as masked vectors (using zero or negative infinity values) to better capture dependencies across available modalities, as illustrated in Figure 9(b). Unlike conventional cross-modal attention, masked attention models share the same parameters across all embeddings, allowing flexible handling of missing modalities. For example, Qian et al. (Qian and Wang, 2023) designed an attention mask matrix to ignore missing modalities, improving model robustness. Similarly, DrFuse(Yao et al., 2024) decouples modalities into specific and shared representations, using the shared ones to replace missing modalities, with a custom mask matrix to help model ignore the specific representations of the missing modality.
@@
Figure 10. Description of general idea of joint representation learning methods, assuming Modality 2 is the missing modality. The missing modality 2 tokens can be replaced by specific masked tokens. The modality encoders (red dashed box) can be replaced by linear projection layers in some methods.
Transformer Based Methods can be divided into two types: joint representation learning (JRL) and parameter efficient learning (PEL) according to full parameter training and a small amount of parameter fine-tuning. Since transformers have long context lengths to handle many feature tokens, the multimodal transformer can learn joint representations from any number of modality tokens (Figure 10). Some methods first use modality encoders to extract feature tokens from available modalities and then feed them into a multimodal transformer. Gong et al. (Gong et al., 2023) introduced an egocentric multimodal task, proposing a transformer-based fusion module with a flexible number of modality tokens and a cross-modal contrast alignment loss to map features into a common space. Similarly, Mordacq et al. (Mordacq et al., 2024) leveraged Masked Multimodal Transformers, treating missing modalities as masked tokens for robust JRL. Ma et al. (Ma et al., 2022) proposed an optimal fusion strategy search strategy within the multimodal transformer to help find the best fusion method for different missing/full-modality representations. Radosavovic et al. (Radosavovic et al., 2024) introduced specially designed mask tokens for an autoregressive transformer to manage missing modalities, successfully deploying the model in real-world scenarios.
@@
With the rise of pre-trained transformer models, PEL methods have been developed to fine-tune these models by training few parameters. Two common PEL methods for pre-trained models are prompt and adapter tuning. Initially used in natural language processing, prompt tuning optimizes input prompts while keeping model parameters fixed. It has since been extended to multimodal models. Jang et al. (Jang et al., 2024) introduced modality-specific prompts to address limitations in earlier methods, which merge when all modalities are present and allow the model to update all learnable prompts during training. Liu et al. (Liu et al., 2024b) further improved this by proposing Fourier Prompt, which uses fast Fourier transform to encode global spectral information of available modalities into learnable prompt tokens which can be used to supplement missing-modality features, enabling cross-attention with feature tokens to address missing modalities. Adapter tuning, on the other hand, involves inserting lightweight adapter layers (e.g., MLPs) into pre-trained models to adapt to new tasks without modifying the original parameters. Qiu et al. (Qiu et al., 2023) proposed a method that used a classifier to identify different missing-modality cases and used intermediate features of that classifier as the missing-modality prompt to cooperate with lightweight Adapters to address missing modality problems.
@@
Although attention-based fusion mechanisms in above methods can effectively help deal with the missing modality problem in any framework, none of them cares about the missing modality that may contain important information required for prediction. In the transformer-based methods, the JRL methods are often limited by a large amount of computing resources and require relatively large datasets to achieve good performance. The PEL methods can achieve efficient fine-tuning, but their performance is still not comparable to that of JRL.
@@
(a) Mean Teacher Distillation Methods
(b) Self Distillation Methods
@@
Figure 12. Description of two process-based distillation methods. (a) “EMA” means exponential moving average. To demonstrate the general architecture of Mean Teacher Distillation, we set the loss on the logits representation. (b) Self-distillation methods for missing modality problems involve the use of models that act as teachers and students, where these models use their own soft labels/representations from each branch to refine themselves during training.
4.1.2. Distillation Based Methods
Knowledge distillation (Hinton et al., 2015) transfers knowledge from a teacher model to a student model. The teacher model, with access to more information, helps the student reconstruct missing modalities. Below, we categorize two types of distillation methods for addressing this problem.
Representation-Based Distillation Methods transfer rich representations from the teacher model to help the student capture and reconstruct missing modality features. We classify them based on whether they use logits or intermediate features. Figure 11 illustrates this type of method. Response distillation methods focus on transferring teacher models’ logits to students, helping it mimic probability distributions. Wang et al. (Wang et al., 2020b) trained modality-specific teachers for missing modalities, then used their soft labels to guide a multimodal student. Hafner and Ban (Hafner and Ban, 2023) employed logits from a teacher model trained with optical data to supervise a reconstruction network for approximating missing optical features from radar data. Pramit et al. (Saha et al., 2024) proposed a Modality-Aware Distillation, leveraging both global and local teacher knowledge in a federated learning setting to help student models to learn how to handle missing modalities.
Intermediate distillation methods align intermediate features between teacher and student models. Shen et al. (Shen and Gao, 2019) used Domain Adversarial Similarity Loss to align intermediate layers of the teacher and student, improving segmentation in missing modality settings. Zhang et al. (Zhang et al., 2023c) applied intermediate distillation in endometriosis diagnosis by distilling features from a TVUS-trained teacher to a student using MRI data.
Process-Based Distillation Methods focus on overall distillation strategies, like Mean Teacher Distillation (MTD) (Tarvainen and Valpola, 2017) and self-distillation(Zhang et al., 2019). These methods emphasize procedural learning over direct representation transfer. MTD enhances stability by using the exponential moving average of the student model’s parameters as a teacher (Figure 12(a)). Chen et al. (Chen et al., 2024a) applied this to missing modality sentiment analysis, treating missing samples as augmented data. Li et al. (Li et al., 2022b) used MTD for Lidar-Radar segmentation, improving robustness against missing modalities.
Self-distillation helps a model improve by learning from its own soft representations (Figure 12(b)). Wang et al. proposed the ShaSpec (Wang et al., 2023a), which utilizes distillation between modality-shared branches and modality-specific branches of available modalities. Based on ShaSpec, researchers proposed Meta-learned Cross-modal Knowledge Distillation (Wang et al., 2024b) to further weigh the importance of different available modalities to improve performance. Ramazanova et al. (Ramazanova et al., 2024) used mutual information and self-distillation for egocentric tasks, making predictions invariant to missing modalities. Shi et al. proposed PASSION (Shi et al., 2024), a self-distillation method designed to use the multi-modal branch to help other uni-modal branches of available modalities to improve multimodal medical segmentation performance with missing modality.
Hybrid Distillation Methods combine various distillation approaches to improve student performance. For medical segmentation, Yang et al. (Yang et al., 2022) distilled teacher model knowledge (including logits and intermediate features) at every decoder layer, outperforming ACNet(Wang et al., 2021). Wang et al. (Wang et al., 2023d) introduced ProtoKD, which captures inter-class feature relationships for improved segmentation under missing modality conditions. Recently, CorrKD (Li et al., 2024a) leverages contrastive distillation and prototype learning to enhance performance in uncertain missing modality cases.
Aforementioned methods address the missing modality problem and achieve good generalization during testing. However, except for some intermediate and self-distillation methods, most assume the complete dataset is available for training (means teacher can access full-modality samples during training), with missing modalities encountered only during testing. Therefore, most distillation methods are unsuitable for handling incomplete training datasets.
@@ 4.1.2. Distillation Based Methods
Figure 13. Illustration of graph fusion methods. We set modality 2 (white circle in the graph) as the missing modalities. Features from modality 1 and 3 can be aggregated via this graph fusion method and fused features maintains the same dimension.
4.1.3. Graph Learning Based Methods
Graph-learning based methods leverage relationships between nodes and edges in graph-structured data for representation learning and prediction. We categorize approaches for addressing the missing modality problem into two main types: graph fusion and graph neural network (GNN) methods.
@@ 4.1.3. Graph Learning Based Methods
Graph Fusion Methods integrate multimodal data using a graph structure (Figure 13), making them adaptable to various networks. For example, Angelou et al. (Angelou et al., 2019) proposed a method mapping each modality to a common space using graph techniques to preserve distances and internal structures. Chen et al. (Chen and Zhang, 2020) introduced HGMF, which builds complex relational networks using hyper edges that dynamically connect available modalities. Zhao et al. (Zhao et al., 2022) developed Modality-Adaptive Feature Interaction for Brain Tumor Segmentation, adjusting feature interactions across modalities based on binary existence codes. More recently, Yang et al. (Yang et al., 2023b) proposed a graph attention based Fusion Block to adaptively fuse multimodal features, using attention-based message passing to share information between modalities. These fusion methods can be flexibly inserted into any network for integrating multiple modalities.
(a) Individual GNN Methods
(b) Unified GNN Methods
@@
Figure 14. Illustration of graph neural network (GNN) based methods. We set modality 2 as the missing modalities. (a) Each modality first passes through DNN or builds its graph, and then they are uniformly processed using GNNs. (b) This method first complete the graph from missing-modality samples via approaches like FeatProp (Malitesta et al., 2024), then process it by GNNs.
@@
Graph Neural Network Methods directly encode multimodal information into a graph structure, with GNNs used to learn and fuse this information. Early approaches (Wang et al., 2015) employed Laplacian graphs to connect complete and incomplete samples. Individual GNN methods (Figure 14(a)), such as DESAlign (Wang et al., 2024c), extract features using neural networks or GNNs and fuse them for prediction. Unified GNN methods ( Figure 14(b)) first complete the graph and then use GNNs for prediction, such as in Zhang et al.’s M3Care (Zhang et al., 2022), which uses adaptive weights to integrate information from similar patients. Lian et al. (Lian et al., 2023) proposed the Graph Completion Network, which reconstructs missing modalities by mapping features back to the input space. In recommendation systems, FeatProp(Malitesta et al., 2024) propagates known multimodal features to infer missing ones, while MUSE (Wu et al., 2024) represents patient-modality relationships as bipartite graphs and learns unified patient representations. In knowledge graphs, Chen et al. (Chen et al., 2023) introduced Entity-Level Modality Alignment to dynamically assign lower weights to missing/uncertain modalities, reducing the risk of misguidance in learning.
The methods above can leverage the graph structure to better capture relationships both within/between modalities and samples. However, these methods tend to have lower efficiency, scalability, and higher development complexity compared to other kind of approaches.
4.1.4. Multimodal Large Language Models (LLM)
The impressive transformative power of LLMs, like ChatGPT (Bubeck et al., 2023), can be explained by their impressive generalization capabilities across many tasks. However, our understanding of the world depends not only on language, but also on other data modalities, like vision and audio. This has led researchers to explore MLLMs, designed to handle diverse user inputs across modalities, including cases with missing modalities, leveraging the flexibility of Transformers. Their architectures are similar to that shown in Figure 10. Due to the idiosyncrasies of MLLMs, we distinguish them from normal transformer-based joint representation learning methods in Section 4.1.1 and introduce them below. In some current MLLM architectures, LLMs act as feature processors, integrating feature tokens from different modality-specific encoders and passing the output to task-/modality-specific decoders. This enables LLM to not only capture rich inter-modal dependencies, but also naturally carry the ability to handle any number of modalities, that is, the ability to solve the missing modality problem.